# Sharp systolic inequalities for invariant tight contact forms on principal $\mathbb{S}^{1}$-bundles over $\mathbb{S}^{2}$

### v1

@@ 3.3 A family of potentials
(Theorem 1.2) We claim it is enough to prove the result for $e=1$. Indeed, $M$ has universal cover $\mathbb{S}^{3}$, and the covering map $\mathbb{S}^{3}\rightarrow M$ has degree $e$. Lifting the $\mathbb{S}^{1}$-action and $\alpha$ by this covering map yields a free action of $\mathbb{S}^{1}$ on $\mathbb{S}^{3}$ of Euler number 1 and an invariant contact form $\tilde{\alpha}\in\Omega(1)$. This contact form has volume $\text{Vol}(\tilde{\alpha})$ equal to $e\text{Vol}(\alpha)$. Moreover, a closed Reeb orbit of $\alpha$ with period $T$ lifts to a closed Reeb orbit of $\tilde{\alpha}$ with period $eT$. In particular, we have $\text{sys}(\tilde{\alpha})=e\hskip 2.84544pt\text{sys}(\alpha)$. Finally, $\alpha$ is Zoll if and only if $\tilde{\alpha}$ is Zoll. The claim follows.
@@
Let $\alpha$ be in $\Omega(1)$. We write $J$ for $J_{i_{0}}$, and $K_{\text{min}}$ for the minimal absolute value of a critical value of $K$. Since $]-K_{\text{min}},K_{\text{min}}[$ is contained in $U_{i_{0}}$ and does not intersect $]K_{i}^{-},K_{i}^{+}[$ when $i\neq i_{0}$, Lemma 3.11 tells us that $J$ is positive on $[-K_{\text{min}},K_{\text{min}}]$. By Lemma 3.8 and 3.11, we have
@@
Additionally, the following holds:
$\begin{split}\int_{-K_{\text{min}}}^{K_{\text{min}}}J\mathrm{d}k&=\int_{-K_{\text{min}}}^{K_{\text{min}}}\left(J+\frac{|k|}{2}\right)\mathrm{d}k-\int_{-K_{\text{min}}}^{K_{\text{min}}}\frac{|k|}{2}\mathrm{d}k\\ &=\int_{-K_{\text{min}}}^{K_{\text{min}}}\left(J+\frac{|k|}{2}\right)\mathrm{d}k-\frac{K_{\text{min}}^{2}}{2}.\end{split}$ (3)
@@
We write $g$ for the function $k\in[-K_{\text{min}},K_{\text{min}}]\mapsto J(k)+\frac{|k|}{2}$. It is continuously differentiable on $]-K_{\text{min}},K_{\text{min}}[$ by virtue of Lemma 3.5 and the assumption $e=1$. We claim that the lower bound
$g(k)\geq\max\left(\frac{|k|}{2},\frac{\text{sys}(\alpha)}{2}\right)$ (4)
holds. Indeed, since $J$ is positive, one has that $g(k)\geq\frac{|k|}{2}$. Furthermore, we have
$g(\pm K_{\text{min}})\geq\frac{K_{\text{min}}}{2}\geq\frac{\text{sys}(\alpha)}{2}.$ (5)
Finally, if $g$ reaches its minimum at a critical point $k_{0}\in]-K_{\text{min}},K_{\text{min}}[$, then:
  • if $k_{0}\neq 0$, we have $|J^{\prime}(k_{0})|=\frac{1}{2}$;
  • if $k_{0}=0$, then for $g$ to be minimal at 0, we must have that $J(k)+\frac{1}{2}k\geq J(0)$ for $k$ positive, implying $J^{\prime}(0^{+})\geq-\frac{1}{2}$. Similarly, we must have $J(k)-\frac{1}{2}k\geq J(0)$ for $k$ negative, which gives us $J^{\prime}(0^{-})\leq\frac{1}{2}$. Since $J^{\prime}(0^{+})-J^{\prime}(0^{-})=-1$, we get that $J^{\prime}(0^{+})=-\frac{1}{2}$ and $J^{\prime}(0^{-})=\frac{1}{2}$.
@@
By Corollary 3.7, $k_{0}$ corresponds to a $2$-periodic point of the first-return map of the surface of section, hence a periodic orbit of the Reeb flow, of period twice the corresponding critical value of $g$. We get that critical values of $g$ are not smaller than $\frac{\text{sys}(\alpha)}{2}$. So, the minimum of $g$ being either reached at ±$\pm K_{\text{min}}$ or at a critical point, we get by (5) that
$\min(g)\geq\frac{\text{sys}(\alpha)}{2}.$
The claim (4) follows. Integrating the lower bound (4) in (3) and substituting it in (2) gives that
$\text{Vol}(\alpha)$ $\geq 2\int_{-K_{\text{min}}}^{K_{\text{min}}}J\mathrm{d}k$
$\geq\int_{-K_{\text{min}}}^{K_{\text{min}}}\max\left(|k|,\text{sys}(\alpha)\right)\mathrm{d}k-K_{\text{min}}^{2}$
$\geq\text{sys}(\alpha)^{2}+K_{\text{min}}^{2}-K_{\text{min}}^{2}$
$\geq\text{sys}(\alpha)^{2}.$
@@
This returns the wanted inequality for $e=1$. Moreover, the equality case is obtained if and only if equality holds in (2) and (4). Equality holds in (3) if and only if $K$ has exactly two critical values $K_{\text{min}}$ and $-K_{\text{min}}$. In particular, all Reeb orbits that do not intersect the surface of section $\Sigma$ are closed, with period $K_{\text{min}}$. Furthermore, equality holds in (4) if and only if $\text{sys}(\alpha)$ equals $K_{\text{min}}$, since otherwise $g$ would not be smooth on $]-K_{\text{min}},K_{\text{min}}[\setminus\{0\}$. This happens if and only if $g$ is constant equal to $\frac{K_{\text{min}}}{2}$, i.e. $J(k)=\frac{K_{\text{min}}}{2}-\frac{|k|}{2}$. In that case, $|J^{\prime}(k)|=\frac{1}{2}$ for all $k$, thus $x_{i_{0},k}$ is a $2$-periodic point of the surface of section. The Reeb orbit starting at any of the $x_{i_{0},k}$ is then closed, with minimal period $2\tau(k)=2(J(k)-kJ^{\prime}(k))=K_{\text{min}}$. We getthat every Reeb orbit of $\alpha$ is closed, with common minimal period $K_{\text{min}}$. In other words, $\alpha$ is Zoll. ∎

### v2

@@ 4 Proof of the main results
We will construct contact forms $\alpha$ in $\Omega(e)$ such that $\frac{\text{sys}(\alpha)^{2}}{\text{Vol}(\alpha)}$ is arbitrarily close to $\frac{1}{2}$. We fix $\eta\in]0,\frac{1}{2+e}[$, $a=\frac{1}{2}-\frac{e\eta}{2}>0$ and $f\colon]-\frac{\eta}{2},a[\rightarrow\mathbb{R}$, a smooth function satisfying:
  • for all $r\in]-\frac{\eta}{2},\frac{\eta}{2}[$, $f(r)=\frac{1}{2}-\frac{e}{2}r$;
  • for all $r\in]\eta,a[$, $f(r)=a-r$;
  • $f$ is convex on $]\frac{\eta}{2},\eta[$.
See Figure 1 for an example of such a function. Taking $f$ to be a convex smoothening of the piecewise linear function defined by
$r\mapsto\begin{cases}\frac{1}{2}-\frac{e}{2}r&\text{ if }-\frac{\eta}{2}<r\leq\frac{e\eta}{e+2}\\ a-r&\text{ if }\frac{e\eta}{e+2}\leq r<a\end{cases}$
gives a function as requested since $\frac{\eta}{2}<\frac{e\eta}{e+2}<\eta$ when $e>2$.
$x$$f(x)$$-\frac{\eta}{2}$$\frac{\eta}{2}$$\eta$$a$$\frac{1}{2}$
Figure 1: Example of a smooth convex function $f$ defined on $]-\frac{\eta}{2},a[$, satisfying $f(r)=\frac{1}{2}-\frac{e}{2}r$ on $]-\frac{\eta}{2},\frac{\eta}{2}[$ and $f(r)=a-r$ on $]\eta,a[$.
We now define two contact forms $\alpha_{+}$ and $\alpha_{-}$ on the solid torus $T=\mathbb{S}^{1}\times D^{2}$ as follows:
$\alpha_{+}=(a-r^{2})\mathrm{d}t+f(a-r^{2})\mathrm{d}s$
and
$\alpha_{-}=-(a-r^{2})\mathrm{d}t-f(a-r^{2})\mathrm{d}s$
where $t$ is the coordinate along the $\mathbb{S}^{1}$ factor and $\left(r,s\right)\in\left[0,\sqrt{a+\frac{\eta}{2}}\right]\times\mathbb{S}^{1}$ are polar coordinates on the 2-disk of radius $\sqrt{a+\frac{\eta}{2}}$. Remark that $\alpha_{+}$ and $\alpha_{-}$ are indeed contact forms since the contact condition $\alpha_{\pm}\wedge\mathrm{d}\alpha_{\pm}>0$ becomes the condition $f(a-r^{2})-(a-r^{2})f^{\prime}(a-r^{2})>0$, which is easily seen to be satisfied. We write $U$ for the neighborhood $\left]\sqrt{a-\frac{\eta}{2}},\sqrt{a+\frac{\eta}{2}}\right]\times\partial T$ of $\partial T$. Let $g\colon U\rightarrow U$ be the diffeomorphism defined by $g(r,s,t)=(2a-r^{2},-(s+et),t)$. Following Remark 3.3, gluing $T$ to itself along $U$ via $f$ produces the manifold $M$ that is the total space of the $\mathbb{S}^{1}$-principal bundle over $\mathbb{S}^{2}$ with Euler number $e$. Furthermore, since
$(g^{*}\alpha_{+})_{r,s,t}=-(a-r^{2})\mathrm{d}t-(f(r^{2}-a)+e(r^{2}-a))\mathrm{d}s$
and since $f(r^{2}-a)+e(r^{2}-a)=f(a-r^{2})$ when $(r^{2}-a)\in]-\frac{\eta}{2},\frac{\eta}{2}[$, we obtain that
$(g^{*}\alpha_{+})_{r,s,t}=-(a-r^{2})\mathrm{d}t-f(a-r^{2})\mathrm{d}s=(\alpha_{-})_{r,s,t}$
holds true on $U$. As a consequence, gluing $\alpha_{+}$ and $\alpha_{-}$ via $g$ produces a contact form $\alpha\in\Omega(e)$. This invariant contact form satisfies the following properties :
  • Its moment map $K^{\alpha}(r,s,t)=\iota_{\partial t}\alpha_{\pm}(r,s,t)=\pm(a-r^{2})$ has image the interval $[-a,a]$, and has exactly two critical points, with critical values ±$\pm a$. They correspond to two closed Reeb orbits that coincide with orbits of the $\mathbb{S}^{1}$-action, which have both period $a$.
  • Since $K^{\alpha}$ only has two critical points, the family of surfaces of sections constructed in Section 3.1 consists in only one surface of section, corresponding to the interval $]-a,a[$ of regular values of $K^{\alpha}$.
  • On $(K^{\alpha}>0)$, the decomposition of $\alpha$ described in Section 3.2 is given by the identity $\alpha_{+}=(a-r^{2})\mathrm{d}t+f(a-r^{2})\mathrm{d}s$. With the notations of Section 3.2, one has $\sigma=\mathrm{d}t$, $K(r,s,t)=a-r^{2}$ and $\beta_{r,s,t}=f(a-r^{2})\mathrm{d}s$. Similarly, one has $\sigma=\mathrm{d}t$, $K(r,s,t)=-(a-r^{2})$ and $\beta_{r,s,t}=-f(a-r^{2})\mathrm{d}s$ on $(K^{\alpha}<0)$.
  • Since there is only one surface of section, we omit the index and write $J^{\alpha}\colon]-a,a[\rightarrow\mathbb{R}$ for the corresponding potential. It follows from the identity $\beta_{r,s,t}=f(a-r^{2})\mathrm{d}s$ on $(K^{\alpha}>0)$ that $J^{\alpha}(k)=f(k)$ for $k\in[0,a[$. Similarly, $J^{\alpha}(k)=f(-k)$ for $k\in]-a,0]$.
Except for the two closed Reeb orbits that coincide with $\mathbb{S}^{1}$-orbits, the closed Reeb orbits of $\alpha$ are in correspondance with points where $(J^{\alpha})^{\prime}$ is rational, with period given by Corollary 3.7.
  • If $|r|>\eta$, then $|(J^{\alpha})^{\prime}(r)|=|f^{\prime}(r)|=1$, and the corresponding closed Reeb orbit has period $a$;
  • if $|r|<\frac{\eta}{2}$, then $|(J^{\alpha})^{\prime}(r)|=|f^{\prime}(r)|=\frac{e}{2}$, and the corresponding closed Reeb orbit has period $\frac{1}{2}$ if $e$ is even, or $1$ if $e$ is odd;
  • if $|r|\in]\frac{\eta}{2},\eta[$ and $|(J^{\alpha})^{\prime}(r)|=|f^{\prime}(r)|=\frac{p}{q}$, then the corresponding closed Reeb orbit has period
    $q(J^{\alpha}(r)-r(J^{\alpha})^{\prime}(r))$ $=q(f(|r|-rf^{\prime}(|r|)))$
    $\geq f(|r|-rf^{\prime}(|r|))$
    $\geq a$
    where the last inequality follows from the convexity of $f$.
As a consequence, $\text{sys}(\alpha)=a$, which converges to $\frac{1}{2}$ when $\eta$ converges to zero. Moreover, Lemma 3.8 implies that $\text{Vol}(\alpha)$ converge to $\frac{1}{2}$ when $\eta$ converges to zero. We conclude that $\lim\limits_{\eta\rightarrow 0}\frac{\text{sys}(\alpha)^{2}}{\text{Vol}(\alpha)}=\frac{1}{2}$, which proves the result. ∎
Remark 4.2.
The contact form $\alpha$ constructed in the proof of Proposition 4.1 is briefly described in [34, Section 3], as a deformation of the round metric on $\mathbb{S}^{2}$ to a spindle orbifold.
We can now prove Theorem 1.2.
Proof.
@@
(Theorem 1.2) Let $\alpha\in\Omega(e)$, where $e>0$. We write $J$ for $J_{i_{0}}$, and $K_{\text{min}}$ for the minimal absolute value of a critical value of $K$. Since $]-K_{\text{min}},K_{\text{min}}[$ is contained in $U_{i_{0}}$ and does not intersect $]K_{i}^{-},K_{i}^{+}[$ when $i\neq i_{0}$, Lemma 3.11 tells us that $J$ is positive on $[-K_{\text{min}},K_{\text{min}}]$. By Lemma 3.8 and 3.11, we have
@@
We first prove that $\text{sys}(\alpha)^{2}\leq\frac{1}{2}\text{Vol}(\alpha)$ when $e\geq 2$. In that case, we write $g_{e}$ for the function $k\mapsto J(k)+|k|$. By virtue of Lemma 3.5, $g_{2}$ is continuously differentiable on $]-K_{\text{min}},K_{\text{min}}[$, and $g_{e}$ is continuously differentiable on $]-K_{\text{min}},K_{\text{min}}[\setminus\{0\}$. We claim that if $e\geq 2$,
$g_{e}(k)\geq\max\left(|k|,\text{sys}(\alpha)\right)$ (3)
Indeed, since $J$ is positive, one has $g_{e}(k)\geq|k|$. Furthermore, as a continuous function, $g_{e}$ reaches its minimum on $[-K_{\text{min}},K_{\text{min}}]$ at a point $k_{0}$. Three cases can occur :
  • If $k_{0}\in\{-K_{\text{min}},K_{\text{min}}\}$, then $g_{e}(k_{0})\geq K_{\text{min}}\geq\text{sys}(\alpha)$;
@@
  • If $k_{0}\in]-K_{\text{min}},K_{\text{min}}[\setminus\{0\}$, then $k_{0}$ is a critical point of $g_{e}$. Thus we have $|J^{\prime}(k_{0})|=1$. By Corollary 3.7, $k_{0}$ corresponds to a fixed point of the first-return map of the surface of section, hence a periodic orbit of the Reeb flow, of period the corresponding critical value of $g_{e}$. In particular, $g_{e}(k_{0})\geq\text{sys}(\alpha)$.
  • If $k_{0}=0$, then for $g_{e}$ to be minimal at 0, we must have that $J(k)+k\geq J(0)$ for $k$ positive, implying $J^{\prime}(0^{+})\geq-1$. Similarly, we must have $J(k)-k\geq J(0)$ for $k$ negative, which gives us $J^{\prime}(0^{-})\leq 1$. Since $J^{\prime}(0^{+})-J^{\prime}(0^{-})=-e$, this can only happen when $e=2$, in which case we have $J^{\prime}(0^{+})=-1$ and $J^{\prime}(0^{-})=1$. As in the previous case, Corollary 3.7 gives that the corresponding Reeb orbit is closed, with period $g_{e}(k_{0})$, which is larger than $\text{sys}(\alpha)$.
We get that $\min(g_{e})\geq\text{sys}(\alpha)$, which together with $g_{e}(k)\geq|k|$ gives the lower bound (3). Additionally, the following holds:
$\begin{split}\int_{-K_{\text{min}}}^{K_{\text{min}}}J\mathrm{d}k&=\int_{-K_{\text{min}}}^{K_{\text{min}}}g_{e}(k)\mathrm{d}k-\int_{-K_{\text{min}}}^{K_{\text{min}}}|k|\mathrm{d}k\\ &=\int_{-K_{\text{min}}}^{K_{\text{min}}}g_{e}(k)\mathrm{d}k-K_{\text{min}}^{2}\\ &\geq\int_{-K_{\text{min}}}^{K_{\text{min}}}\max\left(|k|,\text{sys}(\alpha)\right)\mathrm{d}k-K_{\text{min}}^{2}\\ &=\text{sys}(\alpha)^{2}+K_{\text{min}}^{2}-K_{\text{min}}^{2}\\ &=\text{sys}(\alpha)^{2}.\end{split}$ (4)
The two lower bounds (2) and (4) together yield the systolic inequality for $e\geq 2$, namely that
$\text{sys}(\alpha)^{2}\leq\frac{1}{2}\text{Vol}(\alpha).$ (5)
When $e=1$, we define $g_{1}$ as the function $k\mapsto J(k)+\frac{|k|}{2}$. The lower bound
$g_{1}(k)\geq\max\left(\frac{|k|}{2},\frac{\text{sys}(\alpha)}{2}\right)$ (6)
holds for similar reasons to (3). Furthermore, plugging the lower bound 6 in the equation (2) gives
$\begin{split}\text{Vol}(\alpha)&\geq 2\int_{-K_{\text{min}}}^{K_{\text{min}}}J\mathrm{d}k\\ &=2\int_{-K_{\text{min}}}^{K_{\text{min}}}g_{1}(k)\mathrm{d}k-2\int_{-K_{\text{min}}}^{K_{\text{min}}}|k|\mathrm{d}k\\ &\geq\int_{-K_{\text{min}}}^{K_{\text{min}}}\max\left(|k|,\text{sys}(\alpha)\right)\mathrm{d}k-K_{\text{min}}^{2}\\ &=\text{sys}(\alpha)^{2}+K_{\text{min}}^{2}-K_{\text{min}}^{2}\\ &=\text{sys}(\alpha)^{2}.\end{split}$ (7)
This proves the systolic inequality when $e=1$, namely
$\text{sys}(\alpha)^{2}\leq\text{Vol}(\alpha).$ (8)
@@
We now study the equality case. When $e\geq 2$ (resp. $e=1$), the inequality (5) is an equality if and only equality holds in (2) and (3) (resp. in (2) and (6)). Equality holds in (2) if and only if $K$ has exactly two critical values $K_{\text{min}}$ and $-K_{\text{min}}$. In particular, all Reeb orbits that do not intersect the surface of section $\Sigma$ are closed, with period $K_{\text{min}}$. Furthermore, equality holds in (3) if and only if $J(k)=\text{sys}(\alpha)-|k|=K_{\text{min}}-|k|$, which only happens if $e=2$ by Lemma 3.5. In that case, $|J^{\prime}(k)|=1$ for all $k$, thus $x_{i_{0},k}$ is a fixed point of the surface of section. The Reeb orbit starting at any of the $x_{i_{0},k}$ is then closed, with minimal period $\tau(k)=J(k)-kJ^{\prime}(k)=K_{\text{min}}$.
Similarly, when $e=1$, equality holds in (6) if and only if $J(k)=\frac{\text{sys}(\alpha)}{2}-\frac{|k|}{2}=\frac{K_{\text{min}}}{2}-\frac{|k|}{2}$. In that case, $|J^{\prime}(k)|=\frac{1}{2}$ for all $k$, thus $x_{i_{0},k}$ is a 2-periodic point of the surface of section. The Reeb orbit starting at any of the $x_{i_{0},k}$ is then closed, with minimal period $2\tau(k)=2(J(k)-kJ^{\prime}(k))=K_{\text{min}}$.
In both cases, we get that every Reeb orbit of $\alpha$ is closed, with common minimal period $K_{\text{min}}$. In other words, $\alpha$ is Zoll. It follows from the previous argument that the equality case cannot be reached when $e>2$. Finally, Proposition 4.1 implies that the systolic inequality is sharp also when $e>2$. ∎
4.2 Proof of Corollary 1.5
We define the prime action spectrum of a contact form $\alpha$ on a manifold $M$, denoted by $\sigma_{P}(M,\alpha)$, as the set of prime periods of closed Reeb orbits of $\alpha$. We will make use of the following theorem of Cristofaro-Gardiner and Mazzucchelli (Theorem 1.5 in [19]), stating that two Besse contact forms on a closed three-dimensional manifold are diffeomorphic if and only if they share the same prime action spectrum.
Theorem 4.3 (Theorem 1.5 in [19]).
Let $M$ a closed connected $3$-manifold, and $\alpha_{1},\alpha_{2}$ two Besse contact forms on $M$. Then $\sigma_{P}(M,\alpha_{1})=\sigma_{P}(M,\alpha_{2})$ if and only if there exists a diffeomorphism $\Psi\colon M\rightarrow M$ such that $\Psi^{*}\alpha_{2}=\alpha_{1}$.
We now give the proof of Corollary 1.5.
Proof.
[…]
$\text{sys}_{\text{contr}}(\alpha)^{2}\leq|e|\text{Vol}(\alpha).$
Equality holds if and only if the lifted form $p^{*}\alpha$ is Zoll. Since the contractible systolic ratio $\frac{\text{sys}_{\text{contr}}(\alpha)^{2}}{\text{Vol}(\alpha)}$ does not depend on $p$, whether or not $p^{*}\alpha$ is Zoll does not depend on the precise choice of covering map $p$. If $p^{*}\alpha$ is Zoll, then $\alpha$ is Besse since the Reeb orbits of $p^{*}\alpha$, which are closed, project on $M$ to the Reeb orbits of $\alpha$. However not all Besse forms on $M$ can be lifted to Zoll forms on $\mathbb{S}^{3}$. We spend the rest of the proof on proving a characterization of such Besse forms on $M$.
  • When $e<0$: if $p^{*}\alpha$ is Zoll, it follows for Proposition 2.4 that the lifted circle action on $\mathbb{S}^{3}$ coincides with the Reeb flow of $p^{*}\alpha$. Therefore the free circle action on $M$ coincides with the Reeb flow of $\alpha$, which is hence Zoll. Conversely, the Reeb flow of a Zoll contact form in $\Omega(e)$ coincides with the $\mathbb{S}^{1}$-action, which lifts to a free circle action on $\mathbb{S}^{3}$ via any degree $-e$ covering map. Hence Zoll contact forms in $\Omega(e)$ lift to Zoll contact forms on $\mathbb{S}^{3}$.
  • When $e>0$: if $p^{*}\alpha$ is Zoll, the proof of Theorem 1.2 implies that $K^{p^{*}\alpha}$ has exactly two critical points, meaning that exactly two Reeb orbits of $p^{*}\alpha$ coincide with orbits of the circle action. If $T>0$ is the minimal period of the Reeb flow of $p^{*}\alpha$, those two orbits correspond to two closed Reeb orbits of $\alpha$ of minimal period $\frac{T}{e}$. All the other orbits of $p^{*}\alpha$ project to closed Reeb orbits of $\alpha$ of period $T$ when $e$ is odd, $\frac{T}{2}$ when $e$ is even. Thus if $e\in\{1,2\}$, $\alpha$ is Zoll as well, but otherwise $\alpha$ is only Besse and has two singular orbits, which have the same minimal period. Conversely, Zoll contact forms in $\Omega(e)$ lift to Zoll contact forms on $\mathbb{S}^{3}$ when $e\in\{1,2\}$. Assume now that $e>2$. We claim that if $\alpha$ is a Besse contact form on $M$ with exactly two singular orbits, which have the same minimal period, then $\alpha$ lifts to a Zoll contact form on $\mathbb{S}^{3}$.
    The Reeb flow of any Besse contact form induces a Seifert fibration of $M$, namely a circle action on $M$ with finite isotropy groups. For any coprime integers $k_{1},k_{2}\in\mathbb{Z}\setminus\{0\}$, we define a $\mathbb{S}^{1}$-action on $M$, whose orbits define a Seifert fibration of $M$, as follows : the $\mathbb{S}^{1}$-action on $\mathbb{S}^{3}$ defined by
    $t\cdot(z_{1},z_{2})\in\mathbb{S}^{3}\mapsto(e^{ik_{1}t}z_{1},e^{ik_{2}t}z_{2})$
    commutes with the $\mathbb{Z}/e\mathbb{Z}$ action generated by $(z_{1},z_{2})\mapsto(e^{\frac{2\pi i}{e}}z_{1},e^{-\frac{2\pi i}{e}}z_{2})$. As such it induces a circle action on the quotient $\mathbb{S}^{3}\overset{p_{0}}{\rightarrow}M$ by the latter action. When $(k_{1},k_{2})=(1,-1)$ or $(-1,1)$, this defines a free action on $M$. Moreover, when $(k_{1},k_{2})=(1,1)$ or $(-1,-1)$, the circle action on $\mathbb{S}^{3}$ coincides (up to orientation) with the Reeb flow of the standard contact form $\alpha_{0}=\frac{1}{4i}\sum\limits_{i=1}^{2}\left(\bar{z_{i}}\mathrm{d}z_{i}-z_{i}\mathrm{d}\bar{z_{i}}\right)$, which is Zoll. $\alpha_{0}$ induces a Besse form $\alpha_{1}$ on $M$, with prime action spectrum $\left(\frac{T}{e},T\right)$ if $e$ is odd, $\left(\frac{T}{e},\frac{T}{2}\right)$ if $e$ is even. The circle action on $M$ for $(k_{1},k_{2})=(1,1)$ or $(-1,-1)$ coincides with the Reeb flow of $\alpha_{1}$. Finally, the regular orbits of the circle action on $M$ for $(k_{1},k_{2})=(1,-1)$ or $(-1,1)$ have period $1$ if $e$ is odd, $\frac{1}{2}$ if $e$ is even. It has two singular orbits, which have period $\frac{1}{e}$.
    As a consequence of the classification of Seifert fibrations on lens spaces (see [23, Theorem 5.1]), there exist coprime integers $k_{1},k_{2}\in\mathbb{Z}\setminus\{0\}$ and a constant $T>0$ such that the Reeb flow of $\frac{\alpha}{T}$ is conjugated via a diffeomorphism of $M$ to the $\mathbb{S}^{1}$-action on $M$ associated to $(k_{1},k_{2})$. For the two singular orbits to have the same minimal period, $|k_{1}|$ and $|k_{2}|$ must be equal. Since $k_{1}$ and $k_{2}$ are coprime, they must belong to $\{-1,1\}$. If $k_{1}=-k_{2}$, the induced action on $M$ is free, hence $\alpha$ is Zoll. The proof of Theorem 1.2 implies that this cannot happen when $e>2$. Consequently $k_{1}=k_{2}$, and in that case the prime action spectrum of $\alpha$ is $\left(\frac{T}{e},T\right)$ if $e$ is odd, $\left(\frac{T}{e},\frac{T}{2}\right)$ if $e$ is even. It follows that $\frac{\alpha}{T}$ has the same prime action spectrum as $\alpha_{1}$. Applying Theorem 4.3, we get that there is a diffeomorphism $\psi$ of $M$ such that $\psi^{*}\alpha=\alpha_{1}$. Finally, this gives that $(\psi\circ p_{0})^{*}\alpha=p_{0}^{*}\alpha_{1}$ is a Zoll contact form on $\mathbb{S}^{3}$. As a consequence, when $e>2$ a Besse contact form on $M$ lifts to a Zoll contact form on $\mathbb{S}^{3}$ if and only if it has two singular orbits, which have the same prime period.